one pump can drain a pool in 12 minutes. when a second pump is also used the pool only takes 9 minutes to drain.how long would it take the second pump to drain the pool if it were the only pump
Answer by htmentor(1343)

The pumping rate of the 1st pump = 1 pool per 12 min.
With both pumps working together, the combined pumping rate is 1 pool per 9 min.
So, in 9 minutes, the 1st pump pumped 1/12 * 9 min. = 9/12 = 3/4 of the pool
That means the 2nd pump pumped out 1/4 of the pool in 9 minutes.
So the time required for the 2nd pump to pump out the whole pool is 9*4 = 36 minutes.